# NRFormer: Nationwide Nuclear Radiation Forecasting with Spatio-Temporal Transformer

### v2

@@ Abstract.
nuclear radiation forecasting system, spatio-temporal modeling, imbalance-aware transformer
@@ 1. Introduction
In this paper, we study the national-wide nuclear radiation forecasting problem, which involves collaborative forecasting across distributed monitoring stations throughout the country. Figure 1 shows an illustrative example of radiation forecasting in Japan. The dispersion of nuclear radiation is primarily influenced by the global circulation system (i.e., atmospheric and ocean circulation) (Qiao et al., 2011), manifesting intricate spatio-temporal patterns. Therefore, an effective forecasting model must be capable of comprehending these spatio-temporal correlations from geo-distributed stations. Recent advances in Spatio-Temporal Graph Neural Networks (STGNNs) (Jin et al., 2023; Wu et al., 2020; Shao et al., 2022c; Lan et al., 2022; Zhang et al., 2022; Shao et al., 2022a), which combine Graph Neural Networks (GNNs) with sequential models (e.g., RNN), have shown superiority in various applications such as traffic forecasting (Wu et al., 2019; Bai et al., 2020) and air quality prediction (Han et al., 2023). Nonetheless, existing STGNNs often fall short in capturing the complex dynamics of nuclear radiation, primarily due to the following three major challenges.
@@
Challenge 3: Heterogeneous contextual factors. The variation of radiation levels at each station is also influenced by a variety of contextual factors, including meteorological conditions, terrain variations, and significant events. These factors may interact with each other, jointly affecting the path and intensity of nuclear radiation. How to effectively incorporate heterogeneous contextual factors is imperative for improving the accuracy and reliability of the predictive model.
@@ 3.4. User Interfaces
We developed a web-based real-time forecasting system to provide nationwide radiation level predictions across Japan. The deployed system is presented in Figure 10 and 11 in the Appendix and available at https://NRFormer.github.io.
@@ 4.3. Radiation Propagation Prompting
Table 1. Statistic of radiation and meteorological datasets.
Data Description Japan-4H Japan-1D
Nuclear radiation data # of stations 3,841 3,841
# of timesteps 6,121 1,021
Interval 4 hour 1 day
Time span 3/17/2021 - 1/1/2024
Meteorological data # of stations 235
# of timesteps 24,483
Interval 1 hour
Time span 3/17/2021 - 1/1/2024
@@ 5. Experiments
In this section, we introduce the experiment setup, overall performance, ablation study, and deployed system evaluation. Due to the page limit, we report the details of parameter sensitivity analysis in Appendix A.3.
@@ 5.1.2. Evaluation Metrics
We employed three widely used metrics, including Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE), for model evaluation. Lower values in MAE, RMSE, and MAPE indicate higher forecasting accuracy. Moreover, we follow (Zheng et al., 2015; Yi et al., 2018; Liang et al., 2023) to discuss the errors on predicting sudden changes. We identified the top 10% of samples in the test set with the greatest variation between observed and predicted values as sudden change samples for prediction. We provide detailed information on evaluation metrics in Appendix A.2.
@@ 5.1.3. Baseline
We compare the proposed NRFormer with 11 baseline methods, including Historical Average (HA) (Zhang et al., 2017), XGBoost (Chen and Guestrin, 2016), Linear Regression (LR), DCRNN (Li et al., 2018), Graph WaveNet (Wu et al., 2019), STID (Shao et al., 2022b), DLinear (Zeng et al., 2023), StemGNN (Cao et al., 2020), LightCTS (Lai et al., 2023), PatchTST (Nie et al., 2023) and Koopa (Liu et al., 2024). We provide detailed baselines in Appendix A.1.
5.1.4. Implementation Details
5.2. Overall Performance
5.3. Ablation Study
@@ 6. Related Works
Transformer. The advent of the Transformer architecture (Vaswani et al., 2017) has been making a transformative impact across various fields, particularly in time series forecasting (Zhou et al., 2021; Cao et al., 2023; Chowdhury et al., 2022; Gorbett et al., 2023; Wu et al., 2022a; Wang et al., 2023a; Jia et al., 2024) and graph mining (Zhang et al., 2021; Zhu et al., 2023; Fan et al., 2021; Li et al., 2022; Wu et al., 2022b; Chen et al., 2021; Wang et al., 2023b). Transformer depends on self-attention mechanism to capture data correlations and is capable of modeling complicated and long-range dependencies. Due to the powerful capability, recently considerable endeavors have been made to adapt Transformer to time series and spatio-temporal data (Zha et al., 2021; Feng and Tassiulas, 2022; Liang et al., 2023; Zhang and Yan, 2022; Liu et al., 2023; Li et al., 2023). For instance, by introducing a ProbSparse self-attention mechanism, Informer (Zhou et al., 2021) enhances the Transformer framework for time series forecasting with high efficiency and scalability. As another example, Autoformer(Wu et al., 2021) further pushes the boundaries of time series forecasting by integrating decomposition and auto-correlation mechanism into Transformer. In this paper, we adapt the Transformer model with tailored designs to address the unique challenges in nuclear radiation forecasting.
@@
Figure 7. Ablation study of each contextual feature.
[…]
@@ 7. Conclusion
Appendix A Supplementary of Experiment
A.1. Baselines
In this section, we provide a brief overview of the baseline methods.
  • StemGNN (Cao et al., 2020): StemGNN is a novel framework that captures both inter-series correlations and temporal dependencies within multivariate time-series data in the spectral domain, using a combination of Graph Fourier Transform and Discrete Fourier Transform, thereby enhancing forecasting accuracy without relying on pre-defined priors.
  • LightCTS (Lai et al., 2023): LightCTS is a framework aimed at efficient, lightweight forecasting for correlated time series, balancing accuracy with reduced computational needs. It employs simplified operator stacking and innovative modules for better performance on resource-limited devices.
  • PatchTST (Nie et al., 2023): PatchTST is a Transformer-based model for multivariate time series forecasting, leveraging subseries patching and channel independence to reduce complexity and improve long-term accuracy, also excelling in self-supervised tasks and setting new accuracy benchmarks.
A.2. Evaluation Metrics
We utilize three commonly adopted metrics to evaluate the performance of the model, namely Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). These metrics are defined as follows. MAE, by contrast, measures the average magnitude of errors without considering their direction, offering a clear and straightforward interpretation of overall accuracy:
(12) $MAE=\frac{1}{NK}\sum_{i=1}^{N}\sum_{j=1}^{K}|\hat{\mathbf{y}}_{ij}-\mathbf{y}_{ij}|.$
RMSE calculates the square root of the average squared differences between predicted and observed values, thereby placing greater emphasis on larger errors:
(13) $RMSE=\sqrt{\frac{1}{NK}\sum_{i=1}^{N}\sum_{j=1}^{K}(\hat{\mathbf{y}}_{ij}-\mathbf{y}_{ij})^{2}}.$
Meanwhile, MAPE expresses these errors as a percentage of the actual values, facilitating easier comparisons across different scales and datasets. Together, these metrics provide a comprehensive evaluation of the model’s predictive capabilities and robustness:
(14) $MAPE=\frac{1}{NK}\sum_{i=1}^{N}\sum_{j=1}^{K}|\frac{\hat{\mathbf{y}}_{ij}-\mathbf{y}_{ij}}{\mathbf{y}_{ij}}|,$
where $N$ and $K$ are the number of station and future time steps, respectively. $\hat{\mathbf{y}}_{ij}$ denotes the future prediction and $\mathbf{y}_{ij}$ is the ground truth radiation level of station $i$ at $j$-th time step.
Figure 9. Parameter sensitivity of NRFormer.
[…]
A.3. Parameter Sensitivity Analysis
We evaluate the impact of the number of temporal attention layers $L_{t}$ and the number of spatial attention layers $L_{s}$. The results on MAE, RMSE, and MAPE are reported in Figure 9. First, we vary $L_{t}$ from 1 to 5. As can be seen in Figure 9, the performance shows a fluctuation when increasing $L_{t}$ from 1 to 3, and drops slightly by further increasing $L_{t}$ from 3 to 5 on both Japan-4H and Japan-1D datasets. Overall, we achieve the best performance when setting $L_{t}$ = 3. The possible reason is that a small $L_{t}$ is insufficient to capture temporal correlation information, whereas too large $L_{t}$ may introduce redundant and noisy information for our task, leading to performance degradation. We vary $L_{s}$ from 1 to 5 and observe that our model achieves optimal performance when $L_{s}$ is set to 2. Increasing or decreasing $L_{s}$ beyond this point results in a decline in performance. This is primarily because a lower $L_{s}$ fails to provide adequate spatio-temporal correlation of radiation propagation. Additionally, we noticed a performance drop when more spatial attention layers were added. This phenomenon might be caused by the fact that in an imbalanced spatial distribution scenario, using an excessive number of $L_{s}$ can lead to overfitting, particularly for nodes with highly dense connections.
Figure 10. The deployed web-based radiation monitoring and forecasting system. Left panel: Forecasts for individual radiation monitoring stations, with interactive stations enabling users to access detailed predictions for the coming days. Right panel: Radiation dispersion dynamics across Japan, highlighting spatiotemporal propagation patterns over the next several days.
[…]
Figure 11. The data visualization of nuclear radiation in Japan: the left panel shows the distribution of radiation and meteorological monitoring stations; the top-center chart illustrates the average daily radiation levels across different prefectures; the top-right chart depicts the number of radiation monitoring stations in each prefecture; and the bottom-right chart presents temporal radiation patterns observed across various prefectures.
[…]

### v3

@@ Abstract.
radiation forecasting system, spatio-temporal modeling, imbalance-aware transformer
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737252isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Spatial-temporal systemsccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
In this paper, we study the national-wide nuclear radiation forecasting problem, which involves collaborative forecasting across distributed monitoring stations throughout the country. Figure 1 shows an illustrative example of radiation forecasting in Japan. The dispersion of nuclear radiation is primarily influenced by the global circulation system (i.e., atmospheric and ocean circulation) (Qiao et al., 2011), manifesting intricate spatio-temporal patterns. Therefore, an effective forecasting model must be capable of comprehending these spatio-temporal correlations from geo-distributed stations. Recent advances in Spatio-Temporal Graph Neural Networks (STGNNs) (Jin et al., 2023; Wu et al., 2020; Shao et al., 2022c; Lan et al., 2022; Zhang et al., 2022; Shao et al., 2022a), which combine Graph Neural Networks (GNNs) with sequential models (e.g., RNN), have shown superiority in various applications such as traffic forecasting (Wu et al., 2019; Bai et al., 2020; Lyu et al., 2024; Han et al., 2024) and air quality prediction (Han et al., 2023). Nonetheless, existing STGNNs often fall short in capturing the complex dynamics of nuclear radiation, primarily due to the following three major challenges.
@@
Challenge 3: Heterogeneous contextual factors. The variation of radiation at each station is also influenced by a variety of contextual factors, including meteorological conditions, terrain variations, and significant events. These factors may interact with each other, jointly affecting the path and intensity of nuclear radiation. How to effectively incorporate heterogeneous contextual factors is imperative for improving the model’s accuracy and reliability.
@@ 2. Preliminaries
In this section, we introduce some important definitions and formally define the nuclear radiation forecasting problem.
@@ 3.4. User Interfaces
We developed a web-based real-time forecasting system to provide nationwide radiation level predictions across Japan. The deployed system is presented in Figure 10 and 11 in the Appendix.
@@ 4.3. Radiation Propagation Prompting
Table 2. Statistic of radiation and meteorological datasets.
Data Description Japan-4H Japan-1D
Nuclear radiation data # of stations 3,841 3,841
# of timesteps 6,121 1,021
Interval 4 hour 1 day
Time span 3/17/2021 - 1/1/2024
Meteorological data # of stations 235
# of timesteps 24,483
Interval 1 hour
Time span 3/17/2021 - 1/1/2024
@@ 5. Experiments
In this section, we introduce the experiment setup, overall performance, ablation study, parameter sensitivity analysis, and deployed system evaluation.
@@ 5.1.2. Evaluation Metrics
We employed three widely used metrics, including Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE), for model evaluation. Lower values in MAE, RMSE, and MAPE indicate higher forecasting accuracy. Moreover, we follow (Zheng et al., 2015; Yi et al., 2018; Liang et al., 2023) to discuss the errors on predicting sudden changes. We identified the top 10% of samples in the test set with the greatest variation between observed and predicted values as sudden change samples for prediction.
@@ 5.1.3. Baseline
We compare the proposed NRFormer with 11 baseline methods, including:
  • StemGNN (Cao et al., 2020): StemGNN is a novel framework that captures both inter-series correlations and temporal dependencies within multivariate time-series data in the spectral domain.
  • LightCTS (Lai et al., 2023): LightCTS is a framework aimed at efficient, lightweight forecasting for correlated time series, balancing accuracy with reduced computational needs.
  • PatchTST (Nie et al., 2023): PatchTST is a Transformer-based model for multivariate time series forecasting, leveraging subseries patching and channel independence to reduce complexity and improve long-term accuracy.
@@
5.1.4. Implementation Details
5.2. Overall Performance
5.3. Ablation Study
Figure 7. Parameter sensitivity of different temporal and spatial attention layers.
[…]
5.4. Parameter Sensitivity Analysis
Finally, we study the impacts of the hyper-parameters on the performance of radiation forecasting. We evaluate the impact of the number of temporal attention layers $L_{t}$ and the number of spatial attention layers $L_{s}$. The results on MAE, RMSE, and MAPE are reported in Figure 7. First, we vary $L_{t}$ from 1 to 5. As can be seen in Figure 7, the performance shows a fluctuation when increasing $L_{t}$ from 1 to 3, and drops slightly by further increasing $L_{t}$ from 3 to 5 on both Japan-4H and Japan-1D datasets. Overall, we achieve the best performance when setting $L_{t}$ = 3. The possible reason is that a small $L_{t}$ is insufficient to capture temporal correlation information, whereas too large $L_{t}$ may introduce redundant and noisy information for our task, leading to performance degradation. We vary $L_{s}$ from 1 to 5 and observe that our model achieves optimal performance when $L_{s}$ is set to 2. Increasing or decreasing $L_{s}$ beyond this point results in a decline in performance. This is primarily because a lower $L_{s}$ fails to provide adequate spatio-temporal correlation of radiation propagation. Additionally, we noticed a performance drop when more spatial attention layers were added. This phenomenon might be caused by the fact that in an imbalanced spatial distribution scenario, using an excessive number of $L_{s}$ can lead to overfitting, particularly for nodes with highly dense connections.
Figure 8. Ablation study of each contextual feature.
[…]
@@ 6. Related Works
This section emphasizes related works in two aspects: transformer and spatio-temporal forecasting.
Transformer. The advent of the Transformer architecture (Vaswani et al., 2017) has been making a transformative impact across various fields, particularly in time series forecasting (Zhou et al., 2021; Cao et al., 2023; Chowdhury et al., 2022; Gorbett et al., 2023; Wu et al., 2022a; Wang et al., 2023a; Jia et al., 2024) and graph mining (Zhang et al., 2021; Zhu et al., 2023; Fan et al., 2021; Li et al., 2022; Wu et al., 2022b; Chen et al., 2021; Wang et al., 2023b). The transformer depends on the self-attention mechanism to capture data correlations and is capable of modeling complicated and long-range dependencies. Due to the powerful capability, recently considerable endeavors have been made to adapt Transformer to time series and spatio-temporal data (Zha et al., 2021; Feng and Tassiulas, 2022; Liang et al., 2023; Zhang and Yan, 2022; Liu et al., 2023; Li et al., 2023). For instance, by introducing a ProbSparse self-attention mechanism, Informer (Zhou et al., 2021) enhances the Transformer framework for time series forecasting with high efficiency and scalability. STAEformer (Liu et al., 2023) revolutionizes time series forecasting by adeptly capturing spatio-temporal dynamics, setting new benchmarks in accuracy on real-world datasets with its unique embedding technique. As another example, Autoformer(Wu et al., 2021) further pushes the boundaries of time series forecasting by integrating decomposition and auto-correlation mechanism into Transformer. In this paper, we adapt the Transformer model with tailored designs to address the unique challenges in nuclear radiation forecasting.
@@ 7. Conclusion
Acknowledgements.
This work was supported by the National Key R&D Program of China (Grant No.2023YFF0725004), National Natural Science Foundation of China (Grant No.92370204), the Guangzhou Basic and Applied Basic Research Program under Grant No. 2024A04J3279, Education Bureau of Guangzhou Municipality.